OpenStudy (anonymous):

help plz What is the solution to the system of equations? x+y=5 x-2y=2 (_,_)

OpenStudy (aaronandyson):

x+y=5 --- Eqn 1 x-2y=2--- Eqn 2 first do times 2 eqn 1

OpenStudy (anonymous):

im sorry idk im hrry

OpenStudy (aaronandyson):

thats 2x + 2y = 10

OpenStudy (godlovesme):

https://web.geogebra.org/app/ use the link ^^^ and graph the eqns

OpenStudy (aaronandyson):

so now on subtracting 2x + 2y = 10 x - 2y = 2 2y gets canclled

OpenStudy (aaronandyson):

so x = 8 in that sense

OpenStudy (aaronandyson):

now put x = 8 in either eqn 1 or eqn 2

OpenStudy (anonymous):

8+y=5?

OpenStudy (aaronandyson):

:)

OpenStudy (aaronandyson):

thats 8 - 5 = y

OpenStudy (anonymous):

3?

OpenStudy (aaronandyson):

therefore y = 3

OpenStudy (aaronandyson):

yes 3! :)

OpenStudy (anonymous):

(8,3) is the answer

OpenStudy (anonymous):

right?

OpenStudy (aaronandyson):

yes! :)